INFRARED DIODE LASER SPECTROSCOPY OF JET-COOLED Ni(CO)_{3}(^{13}CO) AND NiCO

\underline{ALONZO MARTINEZ} AND MICHAEL D. MORSE, Department of Chemistry, University of Utah, Salt Lake City, UT 84112.

The two major isotopic species (^{58}Ni and ^{60}Ni) of Ni(CO)_{3}(^{13}CO) and NiCO have been investigated with a recently constructed infrared spectrometer. These molecules were observed by implementing a slit jet pulsed nozzle coupled to a modulated electrical discharge and a lock-in detection system. Typical nozzle pulses were approximately 2ms in duration, during which time the electrical discharge was modulated at a frequency of 15 kHz. Modulation of the electrical discharge allowed the use of a lock-in detector to improve the signal to noise ratio. A lead salt diode laser beam was multipassed across the output of the slit jet for 15 passes using a Perry cell, and the transmitted IR intensity was monitored. Analysis of the resulting spectra for the ^{13}CO and CO stretching modes in Ni(CO)_{3}(^{13}CO) and NiCO respectively, lead to the following constants (1σ errors limits in parentheses):

58Ni(CO)_{3}(^{13}CO) ^{60}Ni(CO)_{3}(^{13}CO)  ^{58}NiCO  ^{60}NiCO
ν_{0} (cm^{−1}) 2022.075753(95) 2021.936884(181) 2010.692843(605) 2010.645425(529)
B′′ (cm^{−1})  0.034736(2)  0.033763(4)  0.151094(8)  0.149598(7)
B′ (cm^{−1})  0.034688(2)  0.033709(3)  0.150244(9)  0.148741(5)